OpenStudy (anonymous):

f(x)=ax^2+2 if x<=2, x+20a+8/x^2+1 if x>2 find a so that f'(2) exists

OpenStudy (anonymous):

Left hand derivative and right hand derivative have to be equal..

OpenStudy (dumbcow):

?? did i misread the function f(x) = x +20a +8/x^2

OpenStudy (dumbcow):

oh i didn't see the "+1"

OpenStudy (dumbcow):

sorry

OpenStudy (dumbcow):

is that one fraction ? please clarify functions with parenthesis or use equation editor

OpenStudy (anonymous):

ok clarified

OpenStudy (dumbcow):

ok to find that derivative use quotient rule \[\rightarrow \frac{f'g - fg'}{g^{2}}\] f = x+20a +8 .................. f' = 1 g = x^2 +1 ..................... g' = 2x \[= \frac{(x^{2}+1)-2x(x+20a+8)}{(x^{2}+1)^{2}}\] \[=\frac{1-x^{2}-40ax-16x}{(x^{2}+1)^{2}}\] Now plug in 2 for x, and set equal to 2ax \[\rightarrow 4a = \frac{1-4-80a-32}{25}\] solve for a from there